# Finding Dino: A plug-and-play framework for unsupervised detection of out-of-distribution objects using prototypes

### v1

@@ Abstract
Detecting and localising unknown or Out-of-distribution (OOD) objects in any scene can be a challenging task in vision. Particularly, in safety-critical cases involving autonomous systems like automated vehicles or trains. Supervised anomaly segmentation or open-world object detection models depend on training on exhaustively annotated datasets for every domain and still struggle in distinguishing between background and OOD objects. In this work, we present a plug-and-play generalised framework - PRototype-based zero-shot OOD detection Without Labels (PROWL). It is an inference-based method that does not require training on the domain dataset and relies on extracting relevant features from self-supervised pre-trained models. PROWL can be easily adapted to detect OOD objects in any operational design domain by specifying a list of known classes from this domain. PROWL, as an unsupervised method, outperforms other supervised methods trained without auxiliary OOD data on the RoadAnomaly and RoadObstacle datasets provided in SegmentMeIfYouCan (SMIYC) benchmark. We also demonstrate its suitability for other domains such as rail and maritime scenes.
Keywords:
OOD object detection Anomaly segmentation Prototype learning1 Introduction
@@
Artificial Intelligence (AI) has become a cornerstone of autonomous systems - especially in the perception of the surroundings. Systems operating in the real world must dynamically adapt to any situation in open-world settings. This means that in any given scene: the system should be able to understand its context, usually through the detection and localisation of relevant objects. To this end, AI models are typically trained extensively based on object categories in the given operational design domain (ODD). With high-quality publicly available datasets, such as Cityscapes [10], RailSem19 [33] and MODD [4], several State-of-the-Art (SOTA) deep neural networks (DNNs) provide outstanding performances for a closed set of object categories. However, they are unable to identify and categorise unknown objects, i.e. objects that do not belong to any of the training classes. One can encounter obstacles that were never learned in the training data like a random animal on the road or unknown floating obstacles in front of unmanned surface vehicles (USVs) in maritime applications. Due to the open world setting, there can be numerous such unknown or out-of-distribution (OOD) objects at any time and it is almost impossible to train DNNs exhaustively with annotated datasets on all possible known object categories and object variations, especially in complex domains such as autonomous driving.
@@
In contrast to image classification, where OOD detection is a well-defined and widely researched topic, the main challenge in the context of camera-based object detection is the explicit distinction between an unknown object and the common background, i.e. anything in the scene that is not relevant. Therefore, most existing approaches rely on supervised anomaly segmentation and the use of selected OOD samples during training [28, 8, 1, 30, 22]. Especially the latter is a severe limitation when trying to detect things that were not known at the training time of a model. Open world object detection [34, 17, 14] on the other hand gained some attention recently but is struggling with the application-dependent understanding what a relevant object is.
@@
In this work, we propose a novel framework for the detection of unknown objects in an image: PRototype-based zero-shot OOD detection Without Labels (PROWL). It can detect an arbitrary number of unknown objects in scenes from any domain in a zero-shot plug-and-play manner, i.e. without any additional model training on the target domain, as illustrated in Fig. 1. It leverages the rich and diverse features from pre-trained foundation models such as DINOv2 [27] to robustly capture the known object categories as prototypes in a prototype feature bank. The similarity to those prototypes is then used to calculate a pixel-level similarity score for each given test image. This score can be thresholded to detect OOD pixels. Additionally, we propose to combine foreground masks provided by unsupervised segmentation with those pixel-level scores to refine them into high-quality masks for individual instances of OOD or unknown objects. Making use of pre-trained visual features from foundation models which provide robust representations of almost any object, PROWL can easily generalize across application domains by simply specifying the list of ODD classes to determine the corresponding prototype feature bank. It can unknown objects without additional model training in just simple inference steps and performs comparably to existing supervised methods. To the best of our knowledge, PROWL is the first end-to-end framework for unsupervised zero-shot OOD detection in any scene without any training in ground truth (GT) classes. Within the framework of PROWL, we compare the performance of several different unsupervised segmentation and detection methods based on the quality of foreground masks generated. In the absence of a direct baseline for unsupervised anomaly segmentation in a multi-object scene, we further compare our results with other supervised segmentation methods from the established SMIYC [7] benchmark based on given metrics and datasets. Further, we show the applicability of PROWL for finding unknowns in diverse scenes like road driving, on railroad tracks, or even maritime.
To summarise, the key contributions of our framework PROWL are:
@@
PROWL is the first unsupervised OOD object detection and segmentation framework that can sufficiently and reliably distinguish OOD objects from background
PROWL is a zero-shot OOD object detection framework that relies on pre-trained features from foundation models without additional training on domain data
@@
PROWL can be applied as an adaptable plug-and-play module generalised to any scene in a new domain without domain-specific training. We demonstrate this by applying to new domains beyond road driving such as rail and maritime domains. For the rail domain, we additionally contribute by creating datasets with inpainted OOD objects
@@ 2 Related Work
In vision tasks, detecting the unknown/OOD object has been formulated under different banners.
@@
Anomaly Detection: Anomaly or Out-of-Distribution (OOD) detection was initially conducted in the context of image classification. OOD detection has been widely used for finding deviations from in-distribution (ID), i.e. training data. It encompasses both deviations in distribution shift such as perturbations, weather, or lighting conditions as well as changes in semantic classes previously not seen during training. Methods originating from image classification focused on developing techniques that aimed to quantify uncertainty in confidence values produced by classification outputs of DNNs (e.g., Maximum Softmax Probability [19], Mahalanobis distance [22]). Other methods find anomalies by estimating the likelihood with generative models [29] or training discriminatively with negative or auxiliary data from OOD samples like ODIN [23], Outlier exposure [20]. Another line of work is finding anomalies or defective parts in an object, such as defective part detection in Industrial Anomaly Detection [25]. Although developed as image-level anomaly / OOD detection, most of these methods can be applied to anomaly segmentation by finding potential anomalies based on the confidences of each pixel.
Anomaly Segmentation: The goal is to predict anomaly probabilities for each pixel in an image. Different works use discriminative or generative approaches [12, 24]. Most methods rely on auxiliary out-of-distribution (OOD) data during training. For example, Max Entropy [8] predicts high entropy in anomalous regions and reduces false positives using a meta-classifier on OOD data. DenseHybrid [15] combines discriminative and generative modeling. However, pixel-based reasoning often produces noisy anomaly scores, especially for border pixels and poorly localized anomalies. Recent approaches focus on mask-based methods that capture anomalies as whole objects. These methods predict regions instead of pixels, resulting in fewer false predictions [9]. RbA [26], EAM [16], Maskomaly [1], and Mask2anomaly [28] utilize mask-based classification. However, all these methods are trained with supervised labels and sometimes with OOD data. Our proposed approach eliminates the need to distinguish between ID and OOD object masks. We learn every object as a foreground mask without the notion of OOD/ID object class. We perform unsupervised foreground segmentation and OOD detection separately.
@@ 3 Method
In this section, we provide an overview of our framework PROWL. As illustrated in the architecture diagram in Fig. 2, PROWL comprises three modules - a plug-and-play prototype matching module (Sec. 3.2), a refinement module that is used to generate foreground masks using SOTA unsupervised segmentation methods (Sec. 3.1) and lastly OOD detection module (Sec. 3.3).
3.1 Preliminaries: Foreground Mask Generation
In the refinement module, we rely on the generation of foreground masks for every object in the image without considering their respective classes. We find that SOTA unsupervised segmentation methods such as STEGO or CutLER can provide such masks with high quality. Let the foreground masks generated using these methods be denoted as . STEGO[18] is an unsupervised semantic segmentation model that distills pre-trained unsupervised visual features from DINO [6] into semantic clusters using contrastive loss, thus discovering and segmenting semantic objects without human supervision for each dataset. CutLER [31] is an approach for training unsupervised object detection and segmentation models. It is trained exclusively on unlabeled ImageNet [11] data without any additional in-domain data. Using pre-trained self-supervised features, it uses MaskCut strategy to discover multiple coarse object masks. Further, it trains a detector through several rounds of self-training to detect multiple foreground objects and corresponding instance segmentation masks. Thus, unlike STEGO, CutLER does not provide dense segmentation as output, rather it provides a zero-shot object detection and instance segmentation for detected foreground objects in the image. Upon varying the detector threshold, CutLER can detect a range of small to large objects.
@@ 3.2 Plug-and-play Prototype Matching Module
The first step in our pipeline is to create the feature bank with prototypes for every class specified in the ODD. After creation of the prototype, every pixel is assigned a class via prototype matching.
Creating the prototype feature bank: We aim to create an offline ‘prototype feature bank’ which are global feature space representation of a specific object class or part of an image. Pre-trained features from foundation models like DINO [6] use knowledge distillation via a teacher-student network for learning in a self-supervised approach. Authors of DINO observe that a self-supervised Vision Transformer (ViT) can learn to a great extent the underlying perceptual grouping of image patches and semantic correspondences across images and image domains. This property is even strengthened for DINOv2 features [27] which are trained on much larger image corpora and distilled to smaller models. Thus, we utilize the robust general-purpose visual features from DINOv2 for every known object class in ODD using a minimal number of samples from the train split. We assume a list of ODD object categories (expert-specified) that one can expect in a given scene. Let it consist of known classes and a list of prototype vectors for each class . Let us assume number of samples contribute to each prototype vector for each class . Let the features predicted from a DINOv2 model () be for an image . For each object instance class, the segmentation mask for the instance is extracted, such that each prototype sample image contains only the given object instance with a zeroed-out background, shown in Fig. 2. The features from the region of interest for the object instance are thus calculated as . Finally, prototype vector for object class is updated as: , i.e. the prototype vector list is extended until object instances of class is added. This process is repeated for all the ODD object classes. Depending on the complexity of the dataset, a number of prototype samples can suffice for each object (see Sec 5.3) as compared to supervised training methods which require large training samples.
Prototype matching: For each input image , the inference output is calculated based on the feature vectors from DINOv2. Now, each pixel is classified into one of the K classes based on cosine similarity between feature vectors of the input image and the prototype vectors , resulting in K heatmaps one for each class. Thus, the prototype heatmap for class is calculated by taking the maximum of cosine similarity between the prototype vectors for respective class and features given as -
The list of prototype heatmaps for all ODD classes is given as, . For each pixel in , the assigned class label () and score () is given as:
@@ 3.3 OOD Detection
The next step consists of the detection of OOD pixels or object masks following the prototype-based classification of every pixel in Eq 2. This is done by comparing the cosine similarity scores obtained in Eq 3 with a given threshold . For every pixel we calculate an inverse normalised cosine similarity score as:
Thus, pixels where is designated as an OOD pixel. We have investigated the per-pixel OOD detection performance directly on the prototype heatmaps obtained in according to Equation 1. Although OOD detection based on DINOv2 features can be quite reliable, the resulting OOD detection tends to be noisy, and often the OOD pixels might not precisely localise the relevant object. Thus, the output of PROWL can be further refined by introducing instance-level foreground masks obtained in Sec. 3.1 in combination with the prototype heatmaps. Since these models were trained on huge generic datasets in a self-supervised manner, they can reliably detect multiple foreground object masks in a scene without the notion of known or unknown object class. These masks tend to capture the objectness of every entity in the scene without learning it as background as in supervised detection cases. Thus, every mask where the majority of pixels is designated as OOD is now considered to be OOD for the entire mask. As shown in Fig. 2, the OOD objects dinosaur and passenger car were correctly detected using prototype heatmaps in PROWL as they did not match with any of the classes listed in the prototype bank, however, additional pixels were also spuriously detected as OOD. While PROWL in combination with foreground masks from CutLER, correctly detected and precisely localised the exact OOD object masks as ‘unknown’.
@@ 4.1 Datasets
We demonstrate the plug-and-play performance of our methods by evaluating them in three different operational design domains (ODD).
Road driving scene - We generated prototypes for the urban driving road ODD scene from Cityscapes [10]. It is a benchmark suite with both pixel and instance-level semantic scene understanding. For evaluation on datasets with real OOD objects in the road scene, we refer to the datasets and benchmark suite for anomaly segmentation in SegmentMeIfYouCan (SMIYC) [7]. It provides two novel real-world datasets: RoadAnomaly21 and RoadObstacle21. RoadAnomaly21 consists of test and validation images with real objects or animals as OOD appearing anywhere in the scene. In contrast, RoadObstacle21 has OOD objects (or obstacles) appearing on the road or ego track. The authors of SMIYC withhold the GT for the test set, where the scores are only accessible by submitting the method to the official benchmark. Further, we also evaluate on FS Static subset of Fishyscapes [2] anomaly segmentation benchmark. It consists of validation images with generic objects taken from PASCAL VOC [13] synthetically overlayed on Cityscapes images.
Rail scene - For rail ODD, we use RailSem19[33] dataset which provides diverse images taken from an ego-perspective of a rail vehicle (trains or trams) along with extensive semantic annotations including rails. Regarding OOD detection, there exist no publicly available datasets with OOD objects in the rail scene yet. Thus, we create our validation data consisting of images. We have used Inpaint Anything [32] which uses a combination of SAM, LaMa, and Stable Diffusion, for inpainting potential OOD objects like animals, cars, carts etc on to the RailSem19 images. More details are given in Suppl.
Maritime scene - Obstacle detection and segmentation in maritime ODD (MODD) scenario is prevalent for autonomous operations of unmanned surface vehicles (USVs). Here, the task is to detect and segment all obstacles beyond sky, sea classes as ‘unknown’. We formulate this as an OOD detection task, where prototypes are generated from train split of MaStr1325 (Maritime Semantic Segmentation Training Dataset) [4]. For OOD detection evaluation, we evaluate on corresponding test dataset. We do not compare with MODS benchmark [5] as we do not include evaluations taking into account multi-modal data with IMU calibrations for shoreline detection.

### v2

# Finding Dino: A Plug-and-Play Framework for Zero-Shot Detection of Out-of-Distribution Objects Using Prototypes

@@ Abstract
Detecting and localising unknown or out-of-distribution (OOD) objects in any scene can be a challenging task in vision, particularly in safety-critical cases involving autonomous systems like automated vehicles or trains. Supervised anomaly segmentation or open-world object detection models depend on training on exhaustively annotated datasets for every domain and still struggle in distinguishing between background and OOD objects. In this work, we present a plug-and-play framework - PRototype-based OOD detection Without Labels (PROWL). It is an inference-based method that does not require training on the domain dataset and relies on extracting relevant features from self-supervised pre-trained models. PROWL can be easily adapted to detect in-domain objects in any operational design domain (ODD) in a zero-shot manner by specifying a list of known classes from this domain. PROWL, as a first zero-shot unsupervised method, achieves state-of-the-art results on the RoadAnomaly and RoadObstacle datasets provided in road driving benchmarks - SegmentMeIfYouCan (SMIYC) and Fishyscapes, as well as comparable performance against existing supervised methods trained without auxiliary OOD data. We also demonstrate its generalisability to other domains such as rail and maritime.
1 Introduction
@@
Artificial Intelligence (AI) has become a cornerstone of autonomous systems - especially in the perception of the surroundings. Systems operating in the real world must dynamically adapt to any situation in open-world settings. This means that in any given scene: the system should be able to understand its context, usually through the detection and localisation of relevant objects. To this end, AI models are typically trained extensively based on closed-set object categories in the given operational design domain (ODD). With high-quality publicly available datasets, such as Cityscapes [10], RailSem19 [35] and MODD [4], several State-of-the-Art (SOTA) deep neural networks (DNNs) provide outstanding performances for a closed set of object categories. However, they are unable to identify and categorise unknown objects, i.e. objects that do not belong to any of the training classes. One can encounter obstacles that were never learned in the training data such as a random animal on the road or unknown floating obstacles in front of unmanned surface vehicles (USVs) in maritime applications. Due to the open world setting, there can be numerous such unknown or out-of-distribution (OOD) objects at any time and it is almost impossible to train DNNs exhaustively with annotated datasets on all possible known object categories and object variations, especially in complex domains such as autonomous driving.
@@
In contrast to image classification, where OOD detection is a well-defined and widely researched topic, the main challenge in the context of camera-based object detection is the explicit distinction between an unknown object and the common background, i.e. anything in the scene that is not relevant. Therefore, most existing approaches rely on supervised anomaly segmentation and often use selected OOD samples during training or fine-tuning as auxiliary OOD data[29, 8, 1, 31, 23]. Especially the latter is a severe limitation when trying to detect things that were not known at the training time of a model. Open world object detection [37, 18, 14] on the other hand gained some attention recently but is struggling with the application-dependent understanding what a relevant object is.
@@
In this work, we propose a novel framework for the detection of unknown objects in an image: PRototype-based OOD detection Without Labels (PROWL). It can detect an arbitrary number of unknown objects in scenes from any domain in a zero-shot plug-and-play manner, i.e. without any additional supervised model training or fine-tuning on samples from the target domain, as illustrated in Fig. 1. It leverages the rich and diverse features from frozen foundation models such as self-supervised trained DINOv2 [28] to robustly capture the known object categories as prototypes in a prototype feature bank. The similarity to those prototypes is then used to calculate a pixel-level similarity score for each given test image. This score can be thresholded to detect OOD pixels. Additionally, we propose to combine foreground masks provided by unsupervised segmentation with those pixel-level scores to refine them into high-quality masks for individual instances of OOD or unknown objects. Making use of pre-trained visual features from foundation models which provide robust representations of almost any object, PROWL can easily generalize across application domains by simply specifying the list of ODD classes to determine the corresponding prototype feature bank. It can detect unknown objects without additional model training in just simple inference steps and performs comparably to existing supervised methods. To the best of our knowledge, PROWL is the first end-to-end framework for zero-shot unsupervised OOD object detection in any scene without any explicit training on ground truth (GT) classes. Within the framework of PROWL, we compare the performance of different unsupervised segmentation and detection methods based on the quality of foreground masks generated. In the absence of a direct baseline for zero-shot anomaly segmentation in a multi-object scene, we further compare our results with other supervised segmentation methods from the SMIYC [7] benchmark based on established metrics and datasets.
In summary, the key contributions of our proposed framework PROWL are:
@@
PROWL is the first zero-shot unsupervised OOD object detection and segmentation framework that can sufficiently and reliably distinguish OOD objects from background.
PROWL relies on frozen features from self-supervised foundation models without need for training or fine-tuning on domain data, by simply creating an offline prototype feature bank using very few object samples per class.
@@
PROWL can be applied as plug-and-play module adaptable to any scene in a new domain without domain-specific training. We demonstrate this by applying to new domains beyond road driving such as rail and maritime domains. For the rail domain, we additionally demonstrate by creating a sample dataset with in-painted OOD objects.
@@ 2 Related Work
In vision tasks, detecting the OOD object has been formulated under different banners.
@@
Anomaly Detection: Anomaly or out-of-distribution (OOD) detection was initially conducted in the context of image classification. OOD detection has been widely used for finding deviations from in-distribution (ID), i.e. training data. It encompasses both deviations in distribution shift such as perturbations, weather, or lighting conditions as well as changes in semantic classes unseen during training. Methods originating from image classification focused on developing techniques that aimed to quantify uncertainty in confidence values produced by classification outputs of DNNs (e.g., Maximum Softmax Probability [20], Mahalanobis distance [23]). Other methods find anomalies by estimating the likelihood with generative models [30] or training discriminatively with negative or auxiliary data from OOD samples like ODIN [24], Outlier exposure [21]. Another line of work is finding anomalies or defective parts in an object, such as defective part detection in Industrial Anomaly Detection [26]. Although developed as image-level anomaly detection, most of these methods can be applied to anomaly segmentation by finding potential anomalies based on the confidences of each pixel.
Anomaly Segmentation: The goal is to predict anomaly probabilities for each pixel in an image. Different works use discriminative or generative approaches [12, 25]. Most methods rely on auxiliary OOD data during training. For example, Max Entropy [8] predicts high entropy in anomalous regions and reduces false positives using a meta-classifier on OOD data. DenseHybrid [16] combines discriminative and generative modeling. However, pixel-based reasoning often produces noisy anomaly scores, especially for border pixels and poorly localized anomalies. Recent approaches focus on mask-based methods that capture anomalies as whole objects. These methods predict regions instead of pixels, resulting in fewer false predictions [9]. RbA [27], EAM [17], Maskomaly [1], Mask2anomaly [29] and S2M [36] utilize mask-based classification. However, all these methods are trained with supervised labels, including additional synthetic data [36] and even sometimes exposed to OOD data (auxiliary data). Our proposed approach learns every possible object in the scene as foreground masks without the notion of OOD/ODD object class and performs OOD object detection separately. It eliminates the need for confidence based anomaly scores from supervised training to discover OOD objects.
@@ 3 Method
In this section, we provide an overview of our framework PROWL. As illustrated in the architecture diagram in Fig. 2, PROWL comprises of three modules: a plug-and-play prototype matching module (Sec. 3.1) followed by OOD detection module (Sec. 3.2), and lastly a refinement module used to generate foreground masks for enhanced OOD detection (Sec. 3.3).
@@ 3.1 Plug-and-play Prototype Matching Module
The first step in our pipeline is to create the feature bank with prototype features for every object class specified in the ODD. Subsequently, every pixel is assigned a class via prototype matching step.
Creating the prototype feature bank: We aim to create an offline ‘prototype feature bank’ which consists of global feature space representations corresponding to domain object classes. Pre-trained features from foundation models like DINO [6] use knowledge distillation via a teacher-student network for learning in a self-supervised approach. Authors[6] observe that a self-supervised Vision Transformer (ViT) can learn to a great extent the underlying perceptual grouping of image patches and semantic correspondences across images and image domains. This property is even strengthened for DINOv2 features [28] which are trained on much larger image corpora and distilled to smaller models. Thus, we utilize the robust general-purpose visual frozen features from such feature extractors for object classes in ODD using a minimal number of samples from the train split. We assume a list of ODD object classes (expert-specified) that one can expect in a given scene, say known classes and a list of prototype vectors for each class . Let us assume object instances contribute to each prototype vector for each class . Let the output of the frozen feature extractor () be for an image . Assuming as embedding dimension, token height and width respectively for given backbone, size of is . For each object instance class, the GT mask is multiplied as a binary mask with to extract instance-specific features. Finally, a spatial averaging on last 2 dimensions is performed to obtain feature representation of a prototype instance, given as:
The prototype vector list is extended with vectors until object instances of class is added, repeating for all the ODD object classes. Depending on the complexity of the dataset, only few prototype samples can suffice for each object (Sec 5.3) as compared to supervised training methods which require lots of training samples.
Prototype matching: For each test image , the inference output of the feature extractor is given as . Using prototype feature bank for K classes, K prototype heatmaps are calculated based on maximum cosine similarity between the prototype vector list and , for respective class given as:
Taking maximum over L instances, heatmaps of size are obtained which are then upsampled to image resolution. The list of prototype heatmaps for all the ODD classes is given as, . For each pixel in , the assigned class label () and score () is given as:
@@ 3.2 OOD Detection
The next step is to detect the OOD pixels following the prototype-based classification of every pixel in Eq 3. This is done by comparing the cosine similarity scores obtained in Eq 4 with a given threshold . For every pixel, we calculate an inverse normalised cosine similarity (INCS) score as:
Thus, pixels where is designated as an OOD pixel, otherwise retains the class label . This per-pixel OOD detection based on prototype heatmaps is usually found to be quite reliable, however it could sometimes show noisy detections when the OOD pixels do not belong to a relevant object. Thus, the output of PROWL can be further refined by introducing instance-level foreground masks given in Sec. 3.3 in combination with the prototype heatmaps.
3.3 Refinement module using Foreground Masks
In the refinement module, we focus on the generation of foreground masks for every object in the image irrespective of their classes. SOTA unsupervised segmentation methods such as STEGO or CutLER can provide such masks with high quality. Let the foreground masks generated using these methods be denoted as . STEGO[19] is a semantic segmentation model that distills pre-trained unsupervised visual features from DINO [6] into semantic clusters using contrastive loss, thus discovering and segmenting semantic objects without human supervision for each dataset. CutLER [33] is an approach for training unsupervised object detection and segmentation models. It is trained exclusively on unlabeled ImageNet [11] data without any additional in-domain data. It uses MaskCut strategy to discover multiple coarse object masks, which is used to train a detector through several rounds of self-training to detect multiple foreground objects and corresponding instance segmentation masks. Contrary to STEGO, CutLER does not provide dense segmentation as output, rather it provides a zero-shot object detection and instance segmentation for detected foreground objects.
Since these models were trained on huge generic datasets in a self-supervised manner, they can reliably detect multiple foreground object masks in a scene without the notion of ODD/OOD object class. These masks tend to capture the objectness of every entity in the scene without learning it as background as in supervised detection cases. Thus, every mask where the majority of pixels is designated as OOD (in Sec. 3.2) is now considered to be OOD for the entire mask. As shown in Figure 2, the OOD objects dinosaur and passenger car were correctly detected using prototype heatmaps in PROWL as they did not have high similarity with any of the ODD classes listed in the prototype bank, however, additional pixels were also spuriously detected as OOD. While PROWL in combination with foreground masks, correctly detected and precisely localised the exact OOD object masks as ‘unknown’.
@@ 4.1 Datasets
We demonstrate the plug-and-play performance of our methods by evaluating them in three different ODD:
Road driving scene - We generated prototypes for the urban driving road ODD scene from Cityscapes [10]. It is a benchmark suite with both pixel and instance-level semantic scene understanding. For evaluation on datasets with real OOD objects in the road scene, we refer to the anomaly segmentation in SegmentMeIfYouCan (SMIYC) [7]. It provides two novel real-world datasets: RoadAnomaly21(RA) and RoadObstacle21(RO). RA consists of test and validation images with real objects or animals as OOD appearing anywhere in the scene. In contrast, RO has OOD objects (or obstacles) appearing on the road or ego track. SMIYC withholds the GT for the test set, where the scores are only accessible by submitting the method to the official benchmark. Further, we also evaluate on FS Static subset of Fishyscapes [2] anomaly segmentation benchmark. It consists of test images with generic objects taken from PASCAL VOC [13] synthetically overlayed on Cityscapes images.
Rail scene - For rail ODD, we use RailSem19[35] dataset which provides diverse images taken from an ego-perspective of a rail vehicle (trains or trams) along with extensive semantic annotations including rails. Regarding OOD detection, there exist no publicly available datasets with OOD objects in the rail scene yet. Thus, we create our in-house in-painted test data as detailed in Suppl. A.
Maritime scene - Obstacle detection and segmentation in maritime ODD (MODD) scenario is prevalent for autonomous operations of unmanned surface vehicles (USVs). Here, the task is to detect and segment all obstacles beyond sky, sea classes as ‘unknown’. We formulate this as an OOD detection task, where prototypes are generated from train split of MaStr1325 (Maritime Semantic Segmentation Training Dataset) [4]. For OOD detection evaluation, we evaluate on corresponding test dataset.